OpenStudy (anonymous):

List the (x)-intercepts and test for symmetry a.)y=x^4-20x^2-125 b.)x^2+4y^2=4

OpenStudy (anonymous):

@jim_thompson5910 Could you help me with this?

jimthompson5910 (jim_thompson5910):

to find the x-intercepts, you plug in y = 0 and solve for x

jimthompson5910 (jim_thompson5910):

y=x^4-20x^2-125 0=x^4-20x^2-125 0 = z^2 - 20z - 125 ... I'm letting z = x^2, so z^2 = x^4 z^2 - 20z - 125 = 0 solve for z to get ???

OpenStudy (anonymous):

wait how do i solve for z?

jimthompson5910 (jim_thompson5910):

have you learned about the quadratic formula?

OpenStudy (anonymous):

ohh yeah I forgot about that

OpenStudy (anonymous):

i'll try it now using it

OpenStudy (anonymous):

I got x-25 and x=-5

OpenStudy (anonymous):

It's z, not x. Reject z =-5 because z = x^2 , it can't be negative, --> just take z = 25 --> x^2 =25 , solve for x